OpenStudy (anonymous):

how to solve this 5x/(7y)-9/x

OpenStudy (anonymous):

\[\frac{ 5x }{ 7y }- \frac{ 9 }{ x } \] ? Are you sure that's the right question? :?

OpenStudy (anonymous):

yes its say combine and simplify

OpenStudy (anonymous):

Ah, okay then. After taking the LCM: \[\frac{ 5x ^{2} - 63y }{ 7xy }\] Is that it? Are there any options?

OpenStudy (anonymous):

idk lol its either right or wrong no choices

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Sure. ^.^